OpenStudy (anonymous):

related rates prob. cylinder fills us with water at a rate of 314ft^3 per min. the radius is 10 ft. how fast is the depth of oil changing (dh/dt)?

OpenStudy (anonymous):

fills up**

OpenStudy (anonymous):

so i guess the volume of the cylinder is \[100\pi h\] right?

OpenStudy (anonymous):

that is \[V=100\pi h\] and so \[V'=100\pi h'\] you are told that \[V'=314\] so solve \[314 = 100 \pi h'\] for \[h'\]

OpenStudy (anonymous):

the volume is not 314, its increasing at a rate of 314 ft^3 per min. which is dv/dt=314

OpenStudy (anonymous):

right. it is \[V'=314\]

OpenStudy (anonymous):

the derivative is 314 that is what you need. the volume is changing

OpenStudy (anonymous):

right, but the way you explained above is finding the height when the volume is changing that fast. i am looking for "how fast" the depth of oil is changing with reference to volume

OpenStudy (anonymous):

in other words, im looking for dh/dt

OpenStudy (anonymous):

yes, that is what i found.

OpenStudy (anonymous):

maybe my notation is confusing. i wrote \[V= 100 \pi h\] and then \[V'= 100\pi h'\]

OpenStudy (anonymous):

you can write \[V= 100 \pi h\] and so \[\frac{dV}{dt}=100 \pi\frac{dh}{dt}\]

OpenStudy (anonymous):

you are given the rate of change of the volume, and asked for the rate of change of the height. i wrote the volume as a function of the height, then differentiated both sides wrt t

OpenStudy (anonymous):

ohhh, okay. thanks a lot. got it

OpenStudy (anonymous):

so your answer should be \[h'=\frac{314}{100 \pi}\] or if you prefer \[\frac{dh}{dt}=\frac{314}{100 \pi}\]

OpenStudy (anonymous):

yw (find it easier to write y' instead of dy/dt, but i realize that the other is most frequently used in related rate problems)